OpenStudy (anonymous):

Simplify the expression, 4xy^20^4(-y)^-3 Just this one question!!!

OpenStudy (anonymous):

4xy^20^4(-y)^-3 is this your equation...? \[\huge 4xy^{(20)^{(\frac{-4}{y^3})}}\]

OpenStudy (anonymous):

no, I'm sorry, the equation is y^2 0^4 (-y)^-3. I should've wrote that better

OpenStudy (anonymous):

y^2 times 0^4 times (-y)^-3 ...?

OpenStudy (anonymous):

Here's a picture of the problem if that helps!

OpenStudy (anonymous):

cool, so 0^4 = 0 (0x0x0x0 = 0) and anything time 0 = 0 so final answer is 0

OpenStudy (anonymous):

all good, make sense?

OpenStudy (anonymous):

make sense! but what do we do with the x and y remaining?

OpenStudy (anonymous):

(4 x 0 x -1) = 0, x y^-1 = x/y^1 = x/y, what do we do with this x/y is what I mean, I get the answer is 0 but I just don't understand what happens to the x and y? @Jack4